### v1

# DiP: Taming Diffusion Models in Pixel Space

Abstract
[…]
Figure 1: Comparison of vanilla latent diffusion model, vanilla pixel diffusion model and our method. Vanilla LDMs utilize VAEs to balance computational efficiency and generation quality. Vanilla pixel diffusion models use small patches to pursue detailed generation quality. Our method achieves high-quality generation while maintaining efficient end-to-end training in pixel space.
1 Introduction
Diffusion models [sohl2015deep, ho2020denoising, song2020score, peebles2023scalable, ramesh2022hierarchical, saharia2022photorealistic, yu2022scaling, xie2024sana, song2020denoising, ho2022classifier, karras2024guiding] have reshaped the landscape of generative visual content. With its outstanding generative capabilities of fidelity and diversity, they have established new state-of-the-art benchmarks across a multitude of tasks, including image synthesis [ci2025describe, chen2023diffusion, ye2023ip, wang2024instantid, zhao2025ultrahr, chen2025ragd, zhou2024migc, zhou20243dis, zhao2024wavelet, chen2023pixart, du2025textcrafter], video generation [fan2025instancecap, nan2024openvid, zhang2021rstnet], and 3D object creation [zhang2024tar3d, zhang2025ar, zhang2024temo], decisively surpassing prior paradigms like Generative Adversarial Networks (GANs) [zhao2024cycle, goodfellow2020generative, radford2015unsupervised, mirza2014conditional, zhu2017unpaired, karras2019style]. However, this generative prowess is underpinned by immense computational demands. Consequently, the inherent trade-off between generation quality and computational efficiency thus stands as one of the most critical challenges in the field of diffusion models today.
To mitigate this challenge, Latent Diffusion Models (LDMs) [rombach2022high] have emerged as the de facto standard. By employing a pre-trained autoencoder (VAE) [kingma2013auto] to compress high-resolution images into a compact latent space, LDMs significantly reduce the computational complexity of the iterative denoising process, as shown in Figure 1(a). Nevertheless, this approach is not without its limitations, including potential information loss [yao2025reconstruction, kilian2024computational, chen2024deep, gupta2024photorealistic] during VAE compression and a non-end-to-end training pipeline.
Figure 2: Our method achieves the best FID score with minimal computational cost. (Note: LDM latency includes VAE. The methods marked with dashed lines ( ) are our estimated latency based on the sampling method in the corresponding paper, and should actually be greater than the marked values. The rest methods are the actual test results in the same hardware environment.)
The most direct solution to eliminate the shortcomings of LDMs is to train a diffusion model in pixel space. However, existing pixel space diffusion models [hoogeboom2024simpler, dhariwal2021diffusion, chen2025pixelflow, kingma2023understanding, hoogeboom2023simple, teng2023relay], particularly those based on the powerful Transformer architecture [vaswani2017attention], face a severe scalability issue. As shown in Figure 1(b), to capture fine-grained details, they typically rely on small input patches (e.g. $\times2$ or $\times4$), causing the input sequence length to grow quadratically with image resolution. This quadratic scaling renders high-resolution training and inference computationally intractable, creating a formidable barrier to their practical application.
In this paper, we aim to resolve this critical trade-off in pixel diffusion models. We propose an efficient pixel space diffusion framework called DiP. As shown in Figure 1(c), for efficient global structure construction, we employ a DiT [peebles2023scalable] backbone. Critically, we configure it to operate on large image patches (e.g., $\times16$). This setting choice drastically reduces the input sequence length, aligning it with that of mainstream LDMs operating in latent space. Consequently, our model achieves computational efficiency comparable to LDMs while remaining entirely VAE-free, enabling it to effectively capture the global layout and semantic content of the image. However, operating on large patches alone inevitably leads to blurry outputs lacking high-frequency details. To address this, we introduce a lightweight Patch Detailer Head (only 0.3% increase in total parameters), which is not a post-processing module but an integral component co-trained with the DiT backbone. For each large patch, it receives contextual features from the DiT and leverages its strong local receptive fields to synthesize the missing high-frequency information. This synergistic design allows the DiT backbone to focus on the computationally demanding task of global consistency, while the efficient Patch Detailer Head specializes in local texture and detail restoration. As demonstrated in Figure 2, our approach sets a new state-of-the-art on the efficiency-quality frontier, achieving superior FID scores at significantly lower latency compared to existing methods. Our main contributions are summarized as follows:
  • We propose DiP, a new end-to-end pixel diffusion model framework that effectively alleviates the trade-off between generation quality and computational efficiency through synergistic global-local modeling.
  • We systematically validate the impact of different architectural designs of our framework, hoping to provide the community with a unified, principled framework.
  • On ImageNet generation benchmarks, our framework achieves state-of-the-art performance and lowest inference latency with low training costs.
2 Related Work
Latent Diffusion Models. Latent Diffusion Models (LDMs) [rombach2022high, podell2023sdxl, esser2024scaling, flux, xie2024sana, chen2023pixart, zhang2023adding] have become the de-facto paradigm for large-scale generative modeling due to their computational efficiency and scalability. By performing the diffusion process in a compressed latent space learned by a VAE, LDMs drastically reduce memory and computational costs, enabling efficient training and inference at high resolutions. Architectural advancements within this paradigm, such as replacing the U-Net [ronneberger2015u] backbone with a more scalable Transformer (DiT) [peebles2023scalable], have further pushed the boundaries of generation quality. Despite their success, this efficiency comes at a cost: the VAE acts as an information bottleneck, imposing a hard ceiling on the final image fidelity and often introducing subtle reconstruction artifacts [skorokhodov2025improving, yao2025reconstruction]. Our work circumvents these limitations by proposing an equally efficient architecture that operates directly in pixel space, thereby eliminating the VAE-induced quality constraints.
Pixel Diffusion Models. Recent years have seen renewed interest in pixel space diffusion models that aim to maximize signal fidelity while addressing computational inefficiency. Early works such as ADM [dhariwal2021diffusion] and DDPM [ho2020denoising] demonstrated the power of diffusion but were constrained by the quadratic complexity of their backbones, rendering them impractical for high resolutions. Multi-scale and image patch-based methods [hoogeboom2024simpler, ding2023patched, chen2025pixelflow, hoogeboom2023simple, hoogeboom2024simpler] further enhance the generation effect by decomposing large images into small image patches. However, these methods essentially simulate locality through brute-force training, which leads to extremely low efficiency. Recent work by PixelNerd [wang2025pixnerd] leverages a Transformer to process image features, which then conditions a NeRF-like coordinate network to act as a renderer for finely reconstructing each image patch, achieving impressive performance. Nevertheless, PixelNerd tightly couples the success of its method with this specific NeRF-like rendering mechanism, which may limit the exploration of a broader design space. We argue that the key to achieving efficient and high-quality pixel space generation lies not in the reliance on a specific structure like NeRF, but rather in the design principle of decoupling global structure construction from local detail refinement. Based on this insight, this paper aims to provide a more principled, efficient, and general solution for pixel space diffusion models.
3 Methods
3.1 Preliminaries
A diffusion process gradually perturbs an initial data sample $\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)$ from the true data distribution into isotropic Gaussian noise:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}}_{t}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\quad\text{ where }\epsilon\sim\mathcal{N}(0,\mathbf{I}),$ (1)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. $\{\beta_{t}\}^{T}_{i=1}$ is a predefined variance schedule that controls the noise level at each step. As $t\rightarrow T,\bar{\alpha}_{t}\rightarrow 0$, and the distribution of $\mathbf{x}_{T}$ converges to a standard normal distribution $p\left(\mathbf{x}_{T}\right)\approx\mathcal{N}(0,\mathbf{I})$.
This discrete formulation can be generalized to a continuous-time setting via a stochastic differential equation (SDE):
$d\mathbf{x}=f(\mathbf{x},t)dt+g(t)d\mathbf{w},$ (2)
where $f(\cdot,t)$ is the drift and $g(t)$ is the diffusion coefficient.
The trajectory of this reverse process is governed by a corresponding probability flow ordinary differential equation (ODE):
$d\mathbf{x}=\left[f(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt.$ (3)
Learning to generate data is thus equivalent to learning the score function $\log p_{t}(\mathbf{x})$ or the associated vector field of this ODE. To train a neural network for this task, several objectives have been proposed. DDPM trains a model $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ to predict the noise component $\theta$ from a noisy sample $x_{t}$:
$\mathcal{L}_{\mathrm{DDPM}}=\mathbb{E}_{t,\mathbf{x}_{0},\epsilon}\left[\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)\right\|^{2}\right].$ (4)
Flow Matching (FM) [esser2024scaling] provides a simulation-free paradigm for directly learning the vector field. It defines a conditional probability path $p_{t}\left(\mathbf{x}\mid\mathbf{x}_{0}\right)$ and a corresponding target vector field $u_{t}(\mathbf{x})$. A network $v_{\theta}(\mathrm{x},t)$ is then trained to regress this field by minimizing the loss:
$\mathcal{L}_{\mathrm{FM}}=\mathbb{E}_{t,p_{t}\left(\mathbf{x}\mid\mathbf{x}_{0}\right)}\left[\left\|u_{t}(\mathbf{x})-v_{\theta}(\mathbf{x},t)\right\|^{2}\right].$ (5)
Figure 3: Overfitting the DiT-only model using a single image in pixel space leads to poor detail reconstruction. Introducing a local inductive bias achieves better reconstruction and accelerates convergence. Please zoom in for details.
Figure 4: The t-SNE visualization of feature space. In the ImageNet validation set, 100 samples were randomly selected from each of the 10 classes for feature visualization. Features are extracted using DiT-only and our method, with each class shown in a distinct color.
Figure 5: Patch Detailer Head with local inductive bias was placed at different locations in the model. The results in Sec. 4.3 show that all three methods offer gains compared to DiT-only.
3.2 Motivation
DiT models the long-range dependencies of an image by partitioning it into a sequence of patches, thereby forming a coherent global structure. However, while the self-attention mechanism excels at modeling macroscopic relationships between patches, it compresses the rich spatial information within each patch into a single, flattened token. This design introduces an inherent limitation: the model can adeptly learn the coarse-level layout and arrangement of patches but struggles to model the fine-grained textures and high-frequency details within each patch, consequently limiting the upper bound of its image generation performance.
To empirically validate this, we conduct a preliminary experiment by overfitting a DiT model on a single high-resolution image in the pixel space. As shown in Figure 3, the model successfully captures the global layout and color palette but fails to render fine textures and sharp edges, resulting in a blurry reconstruction. This result demonstrates that when a DiT architecture operates directly on images, it suffers from a lack of inductive bias [yang2024diffusion, goyal2022inductive, an2024inductive, kadkhodaie2023generalization] at the local level, rendering it incapable of achieving precise pixel-level reconstruction within each patch.
This motivates our core design principle: to augment the global Transformer with a dedicated module that explicitly re-injects this missing inductive bias for local details. In this way, our model can leverage the computational efficiency afforded by large patch sizes while simultaneously generating high-quality images with fine-grained details. As shown in Figure 4, our method achieves tighter intra-class clusters and clearer inter-class separation, whereas vanilla DiT exhibits more mixed distributions. This means that the introduction of local inductive bias can more effectively integrate local textures and edge cues in pixel space and thus improves high-level semantic separability and feature consistency. Such improvements are expected to yield more stable structural alignment and better detail during generation process.
3.3 Framework
Based on the above observation, we introduce a framework for high-quality image generation that operates directly in pixel space. DiP first employs a DiT to model the global structure and long-range dependencies of the image. Subsequently, a lightweight Patch Detailer Head refines the output at the patch level, introducing a local inductive bias to synthesize high-frequency details.
[…]
Local Detail Refinement (Patch Detailer Head). The Patch Detailer Head operates independently and in parallel on each patch. For each patch $i$, it takes two inputs: the corresponding global context map $s_{i}$ and the original noisy pixel patch $p_{i}\in\mathbb{R}^{3\times P\times P}$, where $s_{i}\in\mathbb{R}^{D\times 1\times 1}$ is obtained by reshaping and expanding $S_{\text{global}}$. Its objective is to leverage the global context from $S_{\text{global }}$ to accurately interpret the local noisy information in $p_{i}$, ultimately predicting the corresponding noise component $\epsilon_{i}\in\mathbb{R}^{3\times P\times P}$ for that patch. After processing all $N$ patches in parallel, the resulting sequence of predicted noise patches $\left\{\epsilon_{i}\right\}_{i=1}^{N}$ is reassembled into a full-resolution noise prediction map.
Figure 6: Patch Detailer Head framework. This design introduces the local inductive bias that DiT-only lacks with a low number of parameters, resulting in a high-quality image with rich detail.
Method ImageNet 256×256
[…]
Latent Generative Models
[…]
DiT-XL [peebles2023scalable] 2.27 4.60 278.2 0.83 0.57 2.09s 1400 250x2 675M+86M
MaskDiT-G [zheng2023fast] 2.28 5.67 276.6 0.80 0.61 - 1600 79x2 675M+86M
SiT-XL [ma2024sit] 2.06 4.50 270.3 0.82 0.59 2.09s 1400 250x2 675M+86M
FlowDCN-XL [wang2024flowdcn] 2.00 4.33 263.1 0.82 0.58 - 400 250x2 618M+86M
Pixel Generative Models
[…]
ADM [dhariwal2021diffusion] 3.94 6.14 215.8 0.83 0.53 15.80s 400 500 554M
JetFormer-L [tschannen2024jetformer] 6.64 - - 0.69 0.56 - 500 - 2.8B
SiD [hoogeboom2023simple] 2.77 - 211.8 - - - 800 250×2 2.0B
VDM++ [kingma2023understanding] 2.12 - 278.1 - - - - 250×2 2.46B
RIN [jabri2022scalable] 3.42 - 182.0 - - - 480 1000 410M
Farmer/16 [zheng2025farmer] 3.96 - 250.6 0.79 0.50 - 320 - 1.9B
PixelFlow-XL/4 [chen2025pixelflow] 1.98 5.83 282.1 0.81 0.60 7.50s 320 120x2 677M
DiP-XL/16 2.16 4.79 276.8 0.82 0.61 0.92s 160 100x2 631M
DiP-XL/16 1.98 4.57 282.9 0.80 0.62 0.70s 320 75x2 631M
DiP-XL/16 1.90 4.58 285.4 0.81 0.62 0.92s 320 100x2 631M
Table 1: Comparison of the performance of different methods on ImageNet 256×256 with CFG. Performance metrics are annotated with $\uparrow$ (higher is better) and $\downarrow$ (lower is better). Our method achieves the best FID score. Furthermore, compared to other pixel diffusion models, we achieve the best performance across all metrics with the lowest latency.
3.4 Architecture Design
Exploring Patch Detailer Head Architectures. We investigated several architectures for the Patch Detailer Head, each embodying a different form of inductive bias. Our goal is to present a simple, effective and highly efficient design.
  • Standard MLP. As a simple baseline, we used MLP that takes the feature vector $s_{i}$ and a flattened noisy patch $p_{i}$ as input. While straightforward, this design lacks any inherent spatial bias, treating all pixels within the patch as an unordered set.
  • Coordinate-based MLP. To introduce spatial awareness, a design inspired by NeRF can be adopted [wang2025pixnerd]. For each pixel within $p_{i}$, we concatenate its normalized 2D coordinates. $s_{i}$ is used to dynamically generate the weights of a small, coordinate-based MLP. This implicitly learns a continuous function of the image patch, but it lacks the strong priors for local texture and structure that convolutions provide.
[…]
  • Convolutional U-Net (Our Final Choice). We found that a lightweight convolutional U-Net provided the best performance. The hierarchical structure of downsampling and upsampling paths, combined with skip connections, is exceptionally well-suited for capturing multi-scale spatial features and ensuring local continuity. The inherent inductive biases of convolutions (locality and translation equivariance) are highly effective for denoising local textures and edges. As shown in Figure 6, we instantiate the Patch Detailer Head with a shallow U-Net, which includes 4 downsampling and 4 upsampling blocks. Each block consists of a sequence of Convolution, SiLU activation and pooling layer. The global feature vector $s_{i}$ is concatenated channel-wise with the downsampling output at the bottleneck. This design allows the global semantic information to guide the local refinement process effectively while keeping the parameter count minimal.
We provide experimental evidence for this part in Sec. 4.3. Furthermore, we present a preliminary theoretical analysis in Appendix to model the necessity and effectiveness of the Patch Detailer Head, aiming to offer deeper insights.
Figure 7: Qualitative samples from our model trained at 256 × 256 resolution with classifier-free guidance scale of 4.0. DiP demonstrates fine-grained detail, and high visual quality.
Placement of the Patch Detailer Head. Since the main weakness of DiT backbone being trained directly in pixel space is its lack of local awareness, a natural design question arises: does introducing Patch Detailer Head at different locations in the model also bring gains? As shown in Figure 5, we investigated three placement strategies:
  • Post-hoc Refinement. The Patch Detailer Head is placed only after the final DiT block. This creates a clean separation of concerns: the DiT is solely responsible for global modeling, and Patch Detailer Head is solely responsible for local refinement.
  • Intermediate Injection. The Patch Detailer Head is inserted between DiT blocks. The refined patch representations are then projected back and fed into the subsequent DiT blocks.
  • Hybrid Injection. Patch Detailer Head are placed both at an intermediate stage and at the end of the DiT.
Our experiments in Sec. 4.3 revealed that all three strategies yield comparable performance gains over the baseline DiT. However, the Post-hoc Refinement strategy has a unique advantage: by placing the Head at the end, we treat the standard DiT architecture as a fixed, black-box backbone. This approach requires no modification to the DiT’s internal structure, greatly simplifying implementation and potentially allowing for the use of pre-trained DiT checkpoints. Given its optimal balance between high performance and implementation simplicity, we adopt the post-refinement strategy as the final architecture.
4 Experiments
4.1 Setup
Implementation Details. Our experiments are conducted on the class-conditional ImageNet dataset and original images are center-cropped and resized to $256\times 256$ resolution. We set global batch size to 256. We use DDT [wang2025ddt], a variant of DiT, as our model backbone and apply an Exponential Moving Average (EMA) on the model weights with a decay factor of 0.9999. In Patch Detailer Head, the kernel size of the middle layers is set to 3, the padding to 1, and the kernel size of the last convolutional layer is set to 1. Unless otherwise specified, all samples were generated using the Euler-100 solver. More details are included in Appendix.
Evaluation Protocol. To ensure a comprehensive and rigorous assessment of our model’s generative capabilities, we adhere to the evaluation protocol established by ADM [dhariwal2021diffusion]. We employ a suite of standard quantitative metrics to measure performance across different dimensions. Specifically, we use the Fréchet Inception Distance (FID) [heusel2017gans] to assess overall realism and fidelity, the Spatial FID (sFID) [nash2021generating] to evaluate spatial and structural coherence, and the Inception Score (IS) [salimans2016improved] to measure class-conditional diversity. Furthermore, we report Precision (Prec.)/Recall (Rec.) [kynkaanniemi2019improved] to respectively quantify the fidelity of individual samples and the model’s ability to cover the true data distribution. All metrics are calculated using 50,000 generated samples.
Method ImageNet 256×256
[…]
Scaling Up DiT
[…]
Different Patch Detailer Head
[…]
Table 2: Impact of different design schemes on computational overhead and performance.
4.2 Main Result
Performance. Table 3.3 presents a comprehensive comparison against recent SOTA methods with classifier-free guidance scheduling with guidance interval [kynkaanniemi2024applying]. After 320 training epochs, DiP achieved an FID of 1.90 without requiring a pre-trained VAE, surpassing potentially diffusion models such as DiT-XL (FID 2.27) and SiT-XL (FID 2.06), which require longer training times. DiP outperforms the previous best pixel-based model, PixelFlow-XL/4 (FID 1.98), and significantly exceeds others like ADM (FID 3.94) and VDM++ (FID 2.12). Even with a shorter training schedule of 160 epochs, our model reaches a competitive FID of 2.16, outperforming established models like DiT-XL that require much longer training.
Figure 7 presents qualitative samples of DiP at $256\times 256$ resolution. These visualizations reveal rich detail, demonstrating the effectiveness of introducing local inductive bias. More visualization samples are provided in Appendix.
[…]
4.3 Analysis
In this section, we analyzed the trade-off between generation quality and computational cost during the development of DiP, and at the same time explained the rationality of the Patch Detailer Head we designed.
Patch Detailer Head vs. Scaling Up DiT. A common strategy to improve generative models is to increase the model size. However, our findings indicate that this is a suboptimal approach for pixel space diffusion models. As shown in Table 4.1, increasing the DiT’s depth from 26 to 32 layers yields only a marginal improvement (FID from 5.28 to 4.91) at a considerable cost in parameters and training time. It also means that the effectiveness of our Patch Detailer Head comes from the introduction of effective local inductive biases, rather than increasing network depth.
[…]
Figure 8: The t-SNE visualization of feature space. Features are extracted using Post-hoc Refinement, Intermediate Injection, and Hybrid Injection, with each class shown in a distinct color.
Supplementary Material
Appendix A Why Patch Detailer Head: A Theoretical Perspective
In this section, we try to provide a simplified theoretical analysis to further elucidate why we need local detail refinement in enhancing generation quality. From a general insight, we argue that DiT primarily focuses on the layout and arrangement of the dominant elements in the image, or in other words, the low-frequency signals of the global data. Consequently, it is less effective to learn local details and high-frequency signals. Through the refinement structure, we directly inject all signals from the global data into the learning process, which substantially improves the fine-grained processing of these high-frequency details.
Specifically, we follow the flow matching description of the diffusion process. Given an initial data sample ${\mathbf{x}}_{0}\sim p_{{\rm data}}({\mathbf{x}}_{0})\in\mathbb{R}^{d}$ as the input, a Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I}_{d})$, and $t\in[0,1]$, let
${\mathbf{x}}_{t}=(1-t){\mathbf{x}}_{0}+t\epsilon.$ (6)
In this paper, since all inputs are partitioned into patches of equal size, we first define the patch-level input as follows.
Definition A.1 (Patch-level Input).
For each input ${\mathbf{x}}_{0}\in\mathbb{R}^{d}$, we define the patch-level input as $\left\{{\mathbf{x}}_{0}^{(s)}\right\}_{s=1}^{N}$, where ${\mathbf{x}}_{0}^{(s)}\in\mathbb{R}^{p}$ and ${\mathbf{x}}_{0}=\left[\left({\mathbf{x}}_{0}^{(1)}\right)^{\top},\cdots,\left({\mathbf{x}}_{0}^{(N)}\right)^{\top}\right]^{\top}$, $Np=d$.
It is natural to represent the patch-level input by a series of selection matrices $\left\{\mathbf{P}^{(s)}\right\}_{s=1}^{N}$. For each $s$, $\mathbf{P}^{(s)}\in\mathbb{R}^{p\times d}$ satisfies $\mathbf{P}^{(s)}\left(\mathbf{P}^{(s)}\right)^{\top}=\mathbf{I}_{p}$ and $\mathbf{P}^{(s)}{\mathbf{x}}_{0}={\mathbf{x}}_{0}^{(s)}$. The flow-based models try to minimize a loss function defined as $\mathcal{L}_{{\rm FM}}=\mathbb{E}_{t,{\mathbf{x}}_{0},\epsilon}\left[\left\|f({\mathbf{x}}_{t},t)-(\epsilon-{\mathbf{x}}_{0})\right\|^{2}\right]$. Assuming that each patch is independent of one another, a patch-level predictor $f(\cdot,t)$ tries to estimate the patch-level objective field $\hat{v}^{(s)}=f\left({\mathbf{x}}_{t}^{(s)},t\right)$ for each patch-level noised input ${\mathbf{x}}_{t}^{(s)}=(1-t){\mathbf{x}}_{0}^{(s)}+t\epsilon^{(s)}$, where $\epsilon^{(s)}=\mathbf{P}^{(s)}\epsilon\sim\mathcal{N}(0,\mathbf{I}_{p})$. For the given $\mathcal{L}_{{\rm FM}}$, the optimal predictor is the conditional expectation ^v^(s),* = E[ϵ^(s) -x_0^(s) — x_t^(s) ].
However, in true generation tasks, each patch is not independent of others, because, for natural images, the boundaries between adjacent patches are typically continuous and smoothly varying (e.g., there is little difference between one patch of sky and another). The correlation between patches only weakens when an abrupt transition occurs in the image’s elements, such as at the boundary between sky and grass. Moreover, DiT’s attention-based structure allows a single patch to access partial information from all other patches. Although this information may be coarse, this remains a complex, coupled structure. Therefore, for a DiT model, the estimate of $\hat{v}^{(s)}$ is not only based on ${\mathbf{x}}_{t}^{(s)}$ but also some other information from $\left\{{\mathbf{x}}_{t}^{(l)}\right\}_{l\not=s}$. Thus, we define the effective information below.
Definition A.2 (Effective Information).
For a patch-level noised input $\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}$, we define ${\rm EI}^{(s)}\left(f;\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right)$ to represent the effective information used for a generation model $f$ to estimate the patch-level vector field $\hat{v}^{(s)}$ for any $s\in[N]$.
Assuming that each patch is independent of one another, the patch-level estimate $\hat{v}^{(s)}=f\left({\mathbf{x}}_{t}^{(s)},t\right)$ only uses ${\mathbf{x}}_{t}^{(s)}$ for prediction, which means ${\rm EI}^{(s)}\left(f;\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right)=\left\{{\mathbf{x}}_{t}^{(s)}\right\}$. Thus the optimal predictor can be more generally formulated as $\hat{v}^{(s),*}=\mathbb{E}\left[\epsilon^{(s)}-{\mathbf{x}}_{0}^{(s)}\ \middle|\ {\rm EI}^{(s)}\left(f;\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right)\right]$. For attention-based generation models, we cannot accurately obtain the effective information due to the complex coupling structure. However, based on some standard assumptions on the initial data distribution and some empirical observations, we can still give a brief formulation for the effective information.
Assumption A.3 (Data Distribution).
For the initial data distribution, we assume that $p_{\rm data}\sim\mathcal{N}(\mu,\mathbf{\Sigma})$, where $\mathbf{\Sigma}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, $\mathbf{U}=[{\mathbf{u}}_{1},\cdots,{\mathbf{u}}_{d}]$, and $\mathbf{\Lambda}={\rm diag}\{\lambda_{1},\cdots,\lambda_{d}\}$.
Assumption A.4 (Eigenvalue Decay).
There exists $\alpha>1$ such that for any $i\in[d]$, the eigenvalues of $\mathbf{\Sigma}$ satisfies $\lambda_{i}\asymp i^{-a}$.
Assumption A.3 and A.4 characterize the data distribution as a Gaussian distribution with a covariance of a series of fast-decay eigenvalues. The eigenvalue decay of covariance characterizes the differences in high- and low-frequency signals of the image information. This is consistent with the empirical observation that DiT can effectively learn low-frequency signals but has difficulty capturing high-frequency signals. Given $b>0$, we can decompose the input ${\mathbf{x}}_{0}$ into low- and high-frequency components as
${\mathbf{x}}_{0}=\mu+{\mathbf{x}}_{0,{\rm low}}+{\mathbf{x}}_{0,{\rm high}},$ (7)
where ${\mathbf{x}}_{0,{\rm low}}\sim\mathcal{N}(0,\mathbf{\Sigma}_{{\rm low}})$ and ${\mathbf{x}}_{0,{\rm high}}\sim\mathcal{N}(0,\mathbf{\Sigma}_{{\rm high}})$. $\mathbf{\Sigma}_{{\rm low}}$ satisfies $\mathbf{\Sigma}_{{\rm low}}={\mathbf{U}}_{r}\mathbf{\Lambda}_{r}{\mathbf{U}}_{r}^{\top}=\sum_{i=1}^{r}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}$ where $\lambda_{r}>b$ and $\lambda_{r+1}\leq b$, and $\mathbf{\Sigma}_{{\rm high}}=\mathbf{\Sigma}-\mathbf{\Sigma}_{{\rm low}}$. Thus we can decompose the patch-level noised input ${\mathbf{x}}_{t}^{(s)}$ as
${\mathbf{x}}_{t}^{(s)}=\underbrace{(1-t)\mathbf{P}^{(s)}\mu}_{{\rm Mean}^{(s)}}+\underbrace{(1-t)\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm low}}}_{{\rm Low}^{(s)}}+\underbrace{(1-t)\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm high}}}_{{\rm High}^{(s)}}+\underbrace{t\mathbf{P}^{(s)}\epsilon}_{{\rm Noise}^{(s)}}$ (8)
We can assume that for the DiT model, the effective information is composed of the local patch itself and the low-frequency signals of other patches as below.
Assumption A.5 (EI of DiT).
Given DiT as the predictor, there exists $\beta>0$ such that for any $s\in[N]$, the effective information to estimate the patch-level vector $\hat{v}^{(s)}$ satisfies
${\rm EI}^{(s)}\left({\rm DiT};\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right)=\left\{{\mathbf{x}}_{t}^{(s)}\right\}\cup\left\{{\mathbf{x}}_{t,{\rm low}}^{(l)}\right\}_{l\neq s},$ (9)
where
${\mathbf{x}}_{t,{\rm low}}^{(l)}={\rm Mean}^{(l)}+{\rm Low}^{(l)}+{\rm Noise}^{(l)}$ (10)
for all $l\not=s$.
Our refinement structure directly injects all signals from the initial data ${\mathbf{x}}_{0}$ for prediction, which means that for DiP, the effective information satisfies
${\rm EI}^{(s)}\left({\rm DiP};\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right)={\rm EI}^{(s)}\left({\rm DiT};\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right)\cup\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}=\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}.$ (11)
We define $\hat{v}^{(s)}_{{\rm DiT}}=\mathbb{E}\left[\epsilon^{(s)}-{\mathbf{x}}_{0}^{(s)}\ \middle|\ {\rm EI}^{(s)}\left({\rm DiT};\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right)\right]$ and $\hat{v}^{(s)}_{{\rm DiP}}=\mathbb{E}\left[\epsilon^{(s)}-{\mathbf{x}}_{0}^{(s)}\ \middle|\ {\rm EI}^{(s)}\left({\rm DiP};\left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right)\right]$ as the general near-optimal estimate of DiT and DiP, respectively. Then we obtain the main results below.
Theorem A.6.
Assume that Assumption A.3, A.4 and A.5 hold. Consider using DiT and DiP for the diffusion generation task as the predictor, respectively. The general near-optimal estimate $\hat{v}^{(s)}_{{\rm DiT}}$ and $\hat{v}^{(s)}_{{\rm DiP}}$ satisfy
$\hat{v}^{(s)}_{{\rm DiT}}=\mathbf{P}^{(s)}\hat{\mathbf{B}}\hat{\mathbf{M}}\left({\mathbf{x}}_{t}-(1-t)\mu\right)-\mathbf{P}^{(s)}\mu,$ (12)
and
$\hat{v}^{(s)}_{{\rm DiP}}=\mathbf{P}^{(s)}\mathbf{A}\mathbf{M}\left({\mathbf{x}}_{t}-(1-t)\mu\right)-\mathbf{P}^{(s)}\mu,$ (13)
respectively, where
$\begin{split}&\hat{\mathbf{M}}=\left[(1-t)^{2}\mathbf{\Sigma}_{\rm low}+t^{2}\mathbf{I}_{d}+(1-t)^{2}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\right]^{-1},\\ &\hat{\mathbf{B}}=t\mathbf{I}_{d}-(1-t)\mathbf{B},\qquad\mathbf{B}=\mathbf{\Sigma}_{\rm low}+\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)},\\ &\mathbf{M}=\left[(1-t)^{2}\mathbf{\Sigma}+t^{2}\mathbf{I}_{d}\right]^{-1},\\ &\mathbf{A}=t\mathbf{I}_{d}-(1-t)\mathbf{\Sigma}.\end{split}$ (14)
The denoising operator $\mathbf{P}^{(s)}\hat{\mathbf{B}}\hat{\mathbf{M}}$ and $\mathbf{P}^{(s)}\mathbf{A}\mathbf{M}$ satisfies
$\mathbf{P}^{(s)}\hat{\mathbf{B}}\hat{\mathbf{M}}\asymp\sum_{i=1}^{r}\frac{t-(1-t)\lambda_{i}}{(1-t)^{2}\lambda_{i}+t^{2}}{\mathbf{v}}_{i}{\mathbf{u}}_{i}^{\top}+\sum_{i=r+1}^{d}\frac{\lambda_{i}}{t}{\mathbf{v}}_{i}{\mathbf{u}}_{i}^{\top}+\mathcal{I}_{1}+\mathcal{I}_{2},$ (15)
and
$\mathbf{P}^{(s)}\mathbf{A}\mathbf{M}=\sum_{i=1}^{d}\frac{t-(1-t)\lambda_{i}}{(1-t)^{2}\lambda_{i}+t^{2}}{\mathbf{v}}_{i}{\mathbf{u}}_{i}^{\top},$ (16)
respectively, where ${\mathbf{v}}_{1},\cdots,{\mathbf{v}}_{d}]=\mathbf{P}^{(s)}[{\mathbf{u}}_{1},\cdots,{\mathbf{u}}_{d}]$, $\mathcal{I}_{1}=-\sum_{i=r+1}^{d}\sum_{j=1}^{r}\frac{(1-t)\lambda_{i}}{(1-t)^{2}\lambda_{j}+t^{2}}\left[{\mathbf{u}}_{i}^{\top}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}{\mathbf{u}}_{j}\right]{\mathbf{v}}_{i}{\mathbf{u}}_{j}^{\top}$ and $\mathcal{I}_{2}=-\sum_{i=r+1}^{d}\sum_{j=r+1}^{d}\frac{(1-t)\lambda_{i}}{t^{2}}\left[{\mathbf{u}}_{i}^{\top}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}{\mathbf{u}}_{j}\right]{\mathbf{v}}_{i}{\mathbf{u}}_{j}^{\top}$.
Proof.
See Appendix B. ∎
Remark A.7.
Theorem A.6 illustrates that our designed refinement mechanism exhibits a strong adaptive correction capability for high-frequency signals within the image. Specifically, (12) and (13) align with our objective to estimate the conditional expectation of $\epsilon-{\mathbf{x}}_{0}$ through DiT and DiP. The first term of (12) and (13) represents the estimate of noise, while the second term means the estimate of original data. We focus on the first term regarded as the “denoising process”, with respective “denoising operator” (15) and (16) serve as a global representation characterization, performing denoising on different frequency-domain components derived from the original image.
  • When using only DiT for estimation, Equation (15) indicates that DiT can achieve a good adaptive fit for low-frequency signals ($i\leq r$, dominant signals). However, for high-frequency components in the image, relying solely on DiT may not provide sufficient representational capacity for these components. Specifically,
    • The first term in (15) corresponds to the denoising process applied to the low-frequency signals. Although all these belong to the low-frequency regime, those with larger values of $\lambda_{i}$ ($i\leq r$) contain a stronger proportion of the original image content relative to noise. Consequently, as $\lambda_{i}$ decreases, the denoising operator adaptively selects a larger correction magnitude. (It can be readily demonstrated that $\frac{t-(1-t)\lambda_{i}}{(1-t)^{2}\lambda_{i}+t^{2}}$ exhibits a monotonic increase as $\lambda_{i}$ decreases.)
    • The remaining three terms correspond to the denoising process applied to the high-frequency components. Among them, $\mathcal{I}_{1}$ represents the consistent influence exerted by the low-frequency signal on the high-frequency ones. Since the high-frequency components are associated with $\lambda_{i}$ ($i\geq r+1$) that are significantly smaller than those of the low-frequency regime (Assumption A.4), this term can generally be regarded as $o(\mathbf{1})$. The second term and $\mathcal{I}_{2}$—particularly the second term—point to the following fact: when $t\to 1$ in the early stage of denoising, the magnitude of $\lambda_{i}$ is much smaller than $t$, leading the model to apply only negligible corrections to the high-frequency components. This weakens the model’s ability to learn from this portion of the signal. Conversely, as $t\to 0$ in the late stage of denoising, where the model aims to learn a sensitive compensatory mechanism to capture more fine-grained details from the original image, $t$ becomes much smaller than $\lambda_{i}$, causing the model’s corrections to high-frequency components to lose stability. As a result, these corrections may introduce inconsistencies with the previously learned representations, potentially affecting fine details of the final output.
  • In contrast, when using DiP for estimation, Equation (16) demonstrates that DiP can provide a robust adaptive correction for all signals. This is attributed to our refinement mechanism, which, at a low computational cost, enhances the effective information during the denoising process. Particularly for high-frequency signals, the refinement provides a powerful supplement to the information that DiT struggles to capture, which aligns with both our intuition and experimental results.
Appendix B Proof of Theorem A.6
Proof.
Based on (11), we have $\hat{v}^{(s)}_{{\rm DiP}}=\mathbb{E}\left[\epsilon^{(s)}-{\mathbf{x}}_{0}^{(s)}\ \middle|\ \left\{{\mathbf{x}}_{t}^{(s)}\right\}_{s=1}^{N}\right]=\mathbb{E}\left[\epsilon^{(s)}\ \middle|\ {\mathbf{x}}_{t}\right]-\mathbb{E}\left[{\mathbf{x}}_{0}^{(s)}\ \middle|\ {\mathbf{x}}_{t}\right]$ . We first obtain the following statistics to obtain the first term $\mathbb{E}\left[\epsilon^{(s)}\ \middle|\ {\mathbf{x}}_{t}\right]$. Expectations:
$\begin{split}\mathbb{E}\left[\epsilon^{(s)}\right]=0,\ \ \mathbb{E}\left[{\mathbf{x}}_{t}\right]=(1-t)\mu,\end{split}$ (17)
Covariances:
$\begin{split}\text{Cov}\left(\epsilon^{(s)},{\mathbf{x}}_{t}\right)&=\text{Cov}\left(\mathbf{P}^{(s)}\epsilon,(1-t){\mathbf{x}}_{0}+t\epsilon\right)=t\mathbf{P}^{(s)}\text{Cov}(\epsilon,\epsilon)=t\mathbf{P}^{(s)},\end{split}$ (18)
and
$\begin{split}\text{Cov}({\mathbf{x}}_{t})&=\text{Cov}\left((1-t){\mathbf{x}}_{0}+t\epsilon\right)=(1-t)^{2}\text{Cov}({\mathbf{x}}_{0})+t^{2}\text{Cov}(\epsilon)=(1-t)^{2}\mathbf{\Sigma}+t^{2}\mathbf{I}_{d}.\end{split}$ (19)
Then we use $\mathbb{E}[Y|X]=\mathbb{E}Y+\text{Cov}(Y,X)\text{Cov}(X,X)^{-1}(X-\mathbb{E}X)$ to obtain that
$\mathbb{E}\left[\epsilon^{(s)}\ \middle|\ {\mathbf{x}}_{t}\right]=t\mathbf{P}^{(s)}\left[(1-t)^{2}\mathbf{\Sigma}+t^{2}\mathbf{I}_{d}\right]^{-1}\left({\mathbf{x}}_{t}-(1-t)\mu\right).$ (20)
Similarly, for the second term of $\hat{v}^{(s)}_{{\rm DiP}}$ we have
$\begin{split}\text{Cov}\left({\mathbf{x}}_{0}^{(s)},{\mathbf{x}}_{t}\right)&=\text{Cov}\left(\mathbf{P}^{(s)}{\mathbf{x}}_{0},(1-t){\mathbf{x}}_{0}+t\epsilon\right)=(1-t)\mathbf{P}^{(s)}\text{Cov}({\mathbf{x}}_{0},{\mathbf{x}}_{0})=(1-t)\mathbf{P}^{(s)}\mathbf{\Sigma},\end{split}$ (21)
Then we obtain
$\mathbb{E}\left[{\mathbf{x}}_{0}^{(s)}\ \middle|\ {\mathbf{x}}_{t}\right]=\mathbf{P}^{(s)}\mu+(1-t)\mathbf{P}^{(s)}\mathbf{\Sigma}\left[(1-t)^{2}\mathbf{\Sigma}+t^{2}\mathbf{I}_{d}\right]^{-1}\left({\mathbf{x}}_{t}-(1-t)\mu\right).$ (22)
Thus we have
$\begin{split}\hat{v}^{(s)}_{{\rm DiP}}&=\mathbb{E}\left[\epsilon^{(s)}\ \middle|\ {\mathbf{x}}_{t}\right]-\mathbb{E}\left[{\mathbf{x}}_{0}^{(s)}\ \middle|\ {\mathbf{x}}_{t}\right]\\ &=\mathbf{P}^{(s)}\left[t\mathbf{I}_{d}-(1-t)\mathbf{\Sigma}\right]\left[(1-t)^{2}\mathbf{\Sigma}+t^{2}\mathbf{I}_{d}\right]^{-1}\left({\mathbf{x}}_{t}-(1-t)\mu\right)-\mathbf{P}^{(s)}\mu.\end{split}$ (23)
Letting $\mathbf{M}=\left[(1-t)^{2}\mathbf{\Sigma}+t^{2}\mathbf{I}_{d}\right]^{-1}$, ${\mathbf{v}}_{1},\cdots,{\mathbf{v}}_{d}]=\mathbf{P}^{(s)}[{\mathbf{u}}_{1},\cdots,{\mathbf{u}}_{d}]$, $\mathbf{A}=t\mathbf{I}_{d}-(1-t)\mathbf{\Sigma}$, we have
$\begin{split}\mathbf{P}^{(s)}\mathbf{A}\mathbf{M}&=\left(\sum_{j=1}^{d}{\mathbf{v}}_{j}{\mathbf{u}}_{j}^{\top}\right)\left(\sum_{i=1}^{d}\frac{t-(1-t)\lambda_{i}}{(1-t)^{2}\lambda_{i}+t^{2}}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}\right)\\ &=\sum_{i=1}^{d}\frac{t-(1-t)\lambda_{i}}{(1-t)^{2}\lambda_{i}+t^{2}}{\mathbf{v}}_{i}{\mathbf{u}}_{i}^{\top}.\end{split}$ (24)
Similarly, based on Assumption A.5, we have $\hat{v}^{(s)}_{{\rm DiT}}=\mathbb{E}\left[\epsilon^{(s)}-{\mathbf{x}}_{0}^{(s)}\ \middle|\ \left\{{\mathbf{x}}_{t}^{(s)}\right\}\cup\left\{{\mathbf{x}}_{t,{\rm low}}^{(l)}\right\}_{l\neq s}\right]$, where ${\mathbf{x}}_{t,{\rm low}}^{(l)}=(1-t)\mathbf{P}^{(l)}\mu+(1-t)\mathbf{P}^{(l)}{\mathbf{x}}_{0,{\rm low}}+t\mathbf{P}^{(l)}\epsilon$. We first use one vector to represent the condition $\left\{{\mathbf{x}}_{t}^{(s)}\right\}\cup\left\{{\mathbf{x}}_{t,{\rm low}}^{(l)}\right\}_{l\neq s}$. We try to construct an observation $\hat{\mathbf{x}}_{t}$ such that at $s$ patch, $\mathbf{P}^{(s)}\hat{\mathbf{x}}_{t}={\mathbf{x}}_{t}^{(s)}$, and at $l\not=s$ patch $\mathbf{P}^{(l)}\hat{\mathbf{x}}_{t}={\mathbf{x}}_{t,{\rm low}}^{(l)}$. The following $\hat{\mathbf{x}}_{t}$ satisfies the requirement above
$\hat{\mathbf{x}}_{t}=(1-t)\mu+(1-t){\mathbf{x}}_{0,{\rm low}}+t\epsilon+(1-t)\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm high}}.$ (25)
Now we use the same technique to obtain $\mathbb{E}\left[\epsilon^{(s)}\ \middle|\ \hat{\mathbf{x}}_{t}\right]$. The covariance terms satisfy
$\begin{split}\text{Cov}(\epsilon^{(s)},\hat{\mathbf{x}}_{t})&=\text{Cov}\left(\mathbf{P}^{(s)}\epsilon,t\epsilon\right)=t\mathbf{P}^{(s)},\end{split}$ (26)
and
$\begin{split}\text{Cov}(\hat{\mathbf{x}}_{t})&=\text{Cov}\left((1-t){\mathbf{x}}_{0,{\rm low}}+t\epsilon+(1-t)\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm high}}\right)\\ &=(1-t)^{2}\text{Cov}({\mathbf{x}}_{0,{\rm low}})+t^{2}\text{Cov}(\epsilon)+(1-t)^{2}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\text{Cov}({\mathbf{x}}_{0,{\rm high}})\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\\ &=(1-t)^{2}\mathbf{\Sigma}_{\rm low}+t^{2}\mathbf{I}_{d}+(1-t)^{2}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}.\end{split}$ (27)
Thus we have
$\mathbb{E}\left[\epsilon^{(s)}\ \middle|\ \hat{\mathbf{x}}_{t}\right]=t\mathbf{P}^{(s)}\left[(1-t)^{2}\mathbf{\Sigma}_{\rm low}+t^{2}\mathbf{I}_{d}+(1-t)^{2}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\right]^{-1}\left(\hat{\mathbf{x}}_{t}-(1-t)\mu\right).$ (28)
For $\mathbb{E}\left[{\mathbf{x}}_{0}^{(s)}\ \middle|\ \hat{\mathbf{x}}_{t}\right]$, we have
$\begin{split}\text{Cov}\left({\mathbf{x}}_{0}^{(s)},\hat{\mathbf{x}}_{t}\right)&=\text{Cov}\left(\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm low}}+\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm high}},(1-t)\mu+(1-t){\mathbf{x}}_{0,{\rm low}}+t\epsilon+(1-t)\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm high}}\right)\\ &=\text{Cov}\left(\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm low}},(1-t){\mathbf{x}}_{0,{\rm low}}\right)+\text{Cov}\left(\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm high}},\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}{\mathbf{x}}_{0,{\rm high}}\right)\\ &=(1-t)\mathbf{P}^{(s)}\left[\mathbf{\Sigma}_{\rm low}+\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\right].\end{split}$ (29)
Thus we obtain
$\begin{split}\mathbb{E}\left[{\mathbf{x}}_{0}^{(s)}\ \middle|\ \hat{\mathbf{x}}_{t}\right]=&\mathbf{P}^{(s)}\mu+(1-t)\mathbf{P}^{(s)}\left[\mathbf{\Sigma}_{\rm low}+\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\right]\times\\ &\left[(1-t)^{2}\mathbf{\Sigma}_{\rm low}+t^{2}\mathbf{I}_{d}+(1-t)^{2}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\right]^{-1}\left({\mathbf{x}}_{t}-(1-t)\mu\right).\end{split}$ (30)
Finally we have
$\begin{split}\hat{v}^{(s)}_{{\rm DiT}}&=\mathbb{E}\left[\epsilon^{(s)}\ \middle|\ \hat{\mathbf{x}}_{t}\right]-\mathbb{E}\left[{\mathbf{x}}_{0}^{(s)}\ \middle|\ \hat{\mathbf{x}}_{t}\right]\\ &=\mathbf{P}^{(s)}\left[t\mathbf{I}_{d}-(1-t)\mathbf{B}\right]\hat{\mathbf{M}}\left({\mathbf{x}}_{t}-(1-t)\mu\right)-\mathbf{P}^{(s)}\mu,\end{split}$ (31)
where $\mathbf{B}=\mathbf{\Sigma}_{\rm low}+\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}$ and $\hat{\mathbf{M}}=\left[(1-t)^{2}\mathbf{\Sigma}_{\rm low}+t^{2}\mathbf{I}_{d}+(1-t)^{2}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\mathbf{\Sigma}_{\rm high}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}\right]^{-1}$.
Letting $\hat{\mathbf{B}}=t\mathbf{I}_{d}-(1-t)\mathbf{B}$, we have
$\begin{split}\mathbf{P}^{(s)}\hat{\mathbf{B}}\hat{\mathbf{M}}&=\left(\sum_{j=1}^{d}{\mathbf{v}}_{j}{\mathbf{u}}_{j}^{\top}\right)\left(\hat{\mathbf{C}}_{1}+\hat{\mathbf{C}}_{2}\right)\left(\hat{\mathbf{D}}+\hat{\mathbf{E}}\right)^{-1},\end{split}$ (32)
where
$\begin{split}&\hat{\mathbf{C}}_{1}=\sum_{i=1}^{r}\left(t-(1-t)\lambda_{i}\right){\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}\\ &\hat{\mathbf{C}}_{2}=t\sum_{i=r+1}^{d}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}-(1-t)\sum_{i=r+1}^{d}\lambda_{i}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)},\\ &\hat{\mathbf{D}}=\sum_{i=1}^{r}((1-t)^{2}\lambda_{i}+t^{2}){\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}+\sum_{i=r+1}^{d}t^{2}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top},\\ &\hat{\mathbf{E}}=(1-t)^{2}\sum_{i=r+1}^{d}\lambda_{i}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}{\mathbf{u}}_{i}{\mathbf{u}}_{i}^{\top}\left(\mathbf{P}^{(s)}\right)^{\top}\mathbf{P}^{(s)}.\end{split}$ (33)
We notice that $\hat{\mathbf{D}}$ is a positive diagonal matrix and $\hat{\mathbf{D}}^{-1}\hat{\mathbf{E}}\asymp o(1)$ because Assumption A.4 shows that $\lambda_{p}/\lambda_{q}\asymp(q/p)^{a}\asymp o(1)$ for any $1\leq q\leq r$ and $p\geq r+1$. Thus due to first-order Taylor expansion we have
$\begin{split}\left(\hat{\mathbf{D}}+\hat{\mathbf{E}}\right)^{-1}=\left(\mathbf{I}_{d}+\hat{\mathbf{D}}^{-1}\hat{\mathbf{E}}\right)^{-1}\hat{\mathbf{D}}^{-1}\approx\left(\mathbf{I}_{d}-\hat{\mathbf{D}}^{-1}\hat{\mathbf{E}}\right)\hat{\mathbf{D}}^{-1}=\hat{\mathbf{D}}^{-1}-\hat{\mathbf{D}}^{-1}\hat{\mathbf{E}}\hat{\mathbf{D}}^{-1}\asymp\hat{\mathbf{D}}^{-1}.\end{split}$ (34)
Therefore we obtain
[…]
We finish the proof. ∎
Appendix C More Implementation Details
DiT Architecture
Input dim $\times256$$\times$3
Num. layers 26
Hidden dim. 1152
Num. heads 16
Patch Detailer Head Architecture
[…]
Output Layer 64→3
Optimization
Optimizer AdamW
Learning rate 0.0001
Weight decay 0
Batch size 256
Interpolants
Diffusion sampler Euler
Diffusion steps 100
Evaluation suite ADM
Table 4: Hyperparameter settings.
Hypermarameters.
Table 4 reports the detailed hyperparameters of DiP, including the DiT Architecture, Patch Detailer Head Architecture, Optimization, and Interpolants.
Objective.
DiP follows the training objectives of DDT [wang2025ddt]. It is trained using flow matching as the objective function and regularized using representation alignment techniques. Further improvements could be made by introducing adversarial loss  [goodfellow2014generative], perceptual loss [zhang2018unreasonable].
Sampler.
We use Euler-Maruyama ODE sampler with 100 sampling steps by default. For DiT-only and DiP, we used the same inference hyperparameters.
Classifier-Free Guidance.
In our experiments, we employ Interval-based Classifier-Free Guidance [kynkaanniemi2024applying] (Interval-CFG). Specifically, we set the guidance scale to cfg=2.9. The guidance is activated exclusively within the normalized timestep interval of [0.11, 0.97].
Appendix D How to Preserving High-Frequency Signal: Patch or Image
While the theoretical analysis in Appendix A establishes the need for all-frequency raw signals to refine missing high-frequency details, we also focus on how can this information be injected in the most effective way. Specifically, we are interested in whether patch-level input is better than image-level input, or vise versa, as shown in Figure 11. Intuitively, the transformer structure in DiT has captured the long-distance dependencies, therefore we only need the specific high-frequency signals or details of the image. A toy experiment in Figure 12 verifies this intuition.
In Figure 12(a), through the patch-level input, the learned manifold (black) tightly adheres to the ground truth structure (orange), effectively capturing intricate branching patterns and sharp boundaries. In contrast, Figure 12(b) reveals that global processing leads to over-smoothing. The learned distribution is more dispersed and struggles to lock onto fine structural details. This suggests that we only need the refinement structure to dedicate its capacity to high-frequency sensing without being distracted by long-distance dependencies. On the contrary, with image-level input the network tends to average out features across a broader spatial regime, resulting in a loss of sharp details in high-frequency regions.
Figure 11: Different input formats of Patch Detailer Head.
Figure 12: Toy experiment. (a) Visualization of manifold fitting with Patch-level input. The model precisely captures high-frequency branches. (b) Visualization of manifold fitting with Image-level input. The model exhibits over-smoothing and fails to resolve fine details.
Appendix E More Visualization Results
Figure 13: $\times256$ samples. Class lable = “goldfish, Carassius auratus” (1). CFG = 4.0.
Figure 14: $\times256$ samples. Class lable = “junco, snowbird” (13). CFG = 4.0.
Figure 15: $\times256$ samples. Class lable = “chickadee” (19). CFG = 4.0.
Figure 16: $\times256$ samples. Class lable = “tree frog, tree-frog” (30). CFG = 4.0.
Figure 17: $\times256$ samples. Class lable = “mud turtle” (35). CFG = 4.0.
Figure 18: $\times256$ samples. Class lable = “teddy, teddy bear” (859). CFG = 4.0.
Figure 19: $\times256$ samples. Class lable = “cauliflower” (938). CFG = 4.0.
Figure 20: $\times256$ samples. Class lable = “potpie” (964). CFG = 4.0.
Figure 21: $\times256$ samples. Class lable = “bolete” (997). CFG = 4.0.
Figure 22: $\times512$ samples. Class lable = “ptarmigan” (81). CFG = 4.0.
Figure 23: $\times512$ samples. Class lable=”jellyfish” (107). CFG=4.0.
Figure 24: $\times512$ samples. Class lable=”Maltese dog, Maltese terrier, Maltese” (153). CFG=4.0.
Figure 25: $\times512$ samples. Class lable = “lesser panda, red panda, panda, bear cat, cat bear, Ailurus fulgens” (387). CFG = 4.0.
Figure 26: $\times512$ samples. Class lable = “barn” (425). CFG = 4.0.
Figure 27: $\times512$ samples. Class lable = “beacon, lighthouse, beacon light, pharos” (437). CFG = 4.0.
Figure 28: $\times512$ samples. Class lable = “beer glass” (441). CFG = 4.0.
Figure 29: $\times512$ samples. Class lable = “wool, woolen, woollen” (911). CFG = 4.0.
Figure 30: $\times512$ samples. Class lable = “trifle” (927). CFG = 4.0.

### v2

@@
 Latent Diffusion Models (LDMs) [rombach2022high, podell2023sdxl, esser2024scaling, flux, xie2024sana, chen2023pixart, zhang2023adding] have become the de-facto paradigm for large-scale generative modeling due to their computational efficiency and scalability.
-By performing the diffusion process in a compressed latent space learned by a VAE, LDMs drastically reduce memory and computational costs, enabling efficient training and inference at high resolutions.
+By performing the diffusion process in a compressed latent space learned by a VAE, LDMs drastically reduce memory and computational costs.
 Architectural advancements within this paradigm, such as replacing the U-Net [ronneberger2015u] backbone with a more scalable Transformer (DiT) [peebles2023scalable], have further pushed the boundaries of generation quality.
@@
 Early works such as ADM [dhariwal2021diffusion] and DDPM [ho2020denoising] demonstrated the power of diffusion but were constrained by the quadratic complexity of their backbones, rendering them impractical for high resolutions.
-Multi-scale and image patch-based methods [hoogeboom2024simpler, ding2023patched, chen2025pixelflow, hoogeboom2023simple, hoogeboom2024simpler] further enhance the generation effect by decomposing large images into small image patches.
+Multi-scale and image patch-based methods [hoogeboom2024simpler, ding2023patched, chen2025pixelflow, hoogeboom2023simple, hoogeboom2024simpler] further enhance the generation effect by decomposing large images into small patches.
 However, these methods essentially simulate locality through brute-force training, which leads to extremely low efficiency.
+Concurrent work JiT [li2025jit] demonstrates that high-dimensional data in pixel space can be effectively modeled by predicting clean images.
 Recent work by PixelNerd [wang2025pixnerd] leverages a Transformer to process image features, which then conditions a NeRF-like coordinate network to act as a renderer for finely reconstructing each image patch, achieving impressive performance.
 Nevertheless, PixelNerd tightly couples the success of its method with this specific NeRF-like rendering mechanism, which may limit the exploration of a broader design space.
-We argue that the key to achieving efficient and high-quality pixel space generation lies not in the reliance on a specific structure like NeRF, but rather in the design principle of decoupling global structure construction from local detail refinement.
+We argue that the key to achieving efficient and high-quality pixel space generation lies not in relying on a specific structure like NeRF, but rather in the design principle of decoupling global structure construction from local detail refinement.
 Based on this insight, this paper aims to provide a more principled, efficient, and general solution for pixel space diffusion models.
@@
 DiP-XL/16 1.98 4.57 282.9 0.80 0.62 0.70s 320 75x2 631M
-DiP-XL/16 1.90 4.58 285.4 0.81 0.62 0.92s 320 100x2 631M
-Table 1: Comparison of the performance of different methods on ImageNet 256×256 with CFG.
+DiP-XL/16 1.79 4.59 281.9 0.80 0.63 0.92s 600 100x2 631M
+Table 1: Comparison of the performance of different methods on ImageNet 256×256 with Euler solver and CFG.
 Performance metrics are annotated with $\uparrow$ (higher is better) and $\downarrow$ (lower is better).
@@
 Table 3.3 presents a comprehensive comparison against recent SOTA methods with classifier-free guidance scheduling with guidance interval [kynkaanniemi2024applying].
-After 320 training epochs, DiP achieved an FID of 1.90 without requiring a pre-trained VAE, surpassing potentially diffusion models such as DiT-XL (FID 2.27) and SiT-XL (FID 2.06), which require longer training times.
+After 600 training epochs, DiP achieved an FID of 1.79 without requiring a pre-trained VAE, surpassing potentially diffusion models such as DiT-XL (FID 2.27) and SiT-XL (FID 2.06), which require longer training times.
 DiP outperforms the previous best pixel-based model, PixelFlow-XL/4 (FID 1.98), and significantly exceeds others like ADM (FID 3.94) and VDM++ (FID 2.12).